OpenStudy (anonymous):

Which of the following is a factor of 6x^2 + 2x – 21? explain?

OpenStudy (amistre64):

if it factors nicely ..... multiply the first and last coeffs; what do we get?

OpenStudy (anonymous):

-126?

OpenStudy (amistre64):

and if we create a factor list of -126, we could find 2 factors that add to 2

OpenStudy (anonymous):

none?

OpenStudy (amistre64):

then this wont factor nicely into rational numbers

OpenStudy (anonymous):

so is it prime?

OpenStudy (amistre64):

its not prime, its just not rational ... itll contain a square root in it

OpenStudy (anonymous):

ok so what do i put for an answer?

OpenStudy (amistre64):

6x^2 + 2x – 21 = 0 6x^2 + 2x = 21 6(x^2 + 1/3x) = 21 6(x^2 + 1/3x + 1/36 - 1/36) = 21 6 (x+1/6)^2 - 6/36 = 21 6 (x+1/6)^2 = 21 + 1/6 (x+1/6)^2 = (21(6)+ 1)/36 x+1/6 = +- sqrt(127)/6 x =[-1 +- sqrt(127)]/6

OpenStudy (amistre64):

well considering it says "which of the following"; im not even sure what the options would be ...

OpenStudy (anonymous):

4x – 3 2x + 7 6x – 3 None of the above

OpenStudy (amistre64):

lol, then i know what id pick ....

OpenStudy (amistre64):

it doesnt factor nicely into rationals; and they give you 3 nice rational options

OpenStudy (anonymous):

so is it none of the above

OpenStudy (amistre64):

im going with none of the above, yes

OpenStudy (anonymous):

thank you for explaining and not just giving an answer

OpenStudy (amistre64):

its what i do best ... even when im trying to just give an answer lol